OpenStudy (anonymous):

what is log base 3 of 4

OpenStudy (mathstudent55):

You need to find a value?

OpenStudy (anonymous):

I need to compute it

OpenStudy (mathstudent55):

Ok. Do you have a calculator with logs?

OpenStudy (anonymous):

no I dont

OpenStudy (thecatman):

you can use the calculator on the computer just turn it scientific view i should be there

OpenStudy (mathstudent55):

You can use an online calculator. A calculator will let you calculate logs base e or base 10, but not base 3. That means you need the change of base formula.

OpenStudy (thecatman):

sorry it should be there my key board is malfuntioning

OpenStudy (mathstudent55):

\(\large \log_a x = \dfrac{\log_b x}{\log_b a} \)

OpenStudy (mathstudent55):

You want \(\large \log_3 4\) Using the formula above, you can turn it into a calculation involving log base 10, which a calculator can do. \(\large \log_3 4 = \dfrac{\log_{10} 4}{\log_{10} 3} \)

OpenStudy (anonymous):

so do I do this by hand? or would I punch this formula into a calc

OpenStudy (mathstudent55):

You need a calculator.

OpenStudy (anonymous):

ok, and when I put it in the calculator, i don't need to worry about putting in the 10 right, since that is like the default base?

OpenStudy (mathstudent55):

Correct.

myininaya (myininaya):

your calc (handheld calc) should understand the base is already 10 when nothing is there however some online calc will think the base is something else however it doesn't matter in this case you could do \[\frac{\log_{anything}(4)}{\log_{anything}(3)}\] those anythings have to be same as each other of course and the anything has to be a positive number not equal to 1

OpenStudy (anonymous):

ok, I got 1.262 (rounded third decimal place) is this correct

OpenStudy (mathstudent55):

Correct. That's what I got.

OpenStudy (mathstudent55):

Now if you want to check, raise 3 to that power. You should get 4. That's what logs are all about.

OpenStudy (mathstudent55):

BTW, your calculator probably also has log base e, which is called the natural log, or ln, and that would have worked too. The change of base formula allows you to change to any base you want. Changing to base 10 or base e makes sense bec calculators usually can calculate those two bases easily with just a press of a button.